5. Using your answers to Problems 4a and 4b, estimate your portfolio’s expected return if the security market line is estimated as E(Ri) = 5.2 + 8.4(?i).

Beta of a portfolio is the weighted average Beta of the stock contained in it. Thus,

Problems 4a and 4b answer

1. Beta = (0.23 + 0.77 + 1.05 + 1.33) x 0.25 = 0.845 or 0.85

2. Beta = (0.23 x 20%) + (0.77 x 40%) + (1.05 x 15%) + (1.33 x 25%) = 0.844 or 0.84
Answer #1

If Beta is 0.85 then E(Ri) :-

E(Ri) = 5.2 + 8.4 (0.85) = 12.34%

If Beta is 0.84 then E(Ri) :-

E(Ri) = 5.2 + 8.4 (0.84) = 12.26%